OpenStudy (this):

Solve: 6x2 – 15x = 0

OpenStudy (anonymous):

Take 3x common : 3x(2x - 5) = 0 Either 3x = 0 Or 2x - 5 = 0 So either x = 0 or x = 5/2...

OpenStudy (this):

@waterineyes thanks a lot u got the right answer :D